C(x) = 130x/100-x , 0 ≤ x < 100
describes the cost, C, in millions of dollars, to inoculate x% of the population against a particular strain of the flu. Determine the difference in cost between inoculating 70% of the population and inoculating 40% of the population. (Round to the nearest tenth, if necessary.)

Answer by Theo(13342)

cost to inoculate is (130 * x) / (100 - x)

x is the percent of the population that is being inoculated.

the cost to inoculate 70% of the population would be (130 * 70) / (100 - 70) = (130 * 70) / 30.

this is equal to 303.333333 million dollars

the cost to inoculate 40% of the population would be (130 * 40) / (100 - 40) = (130 * 40) / 60

this is equal to 86.666667 million dollars

the difference is 303.333333 - 86.666667 = 216.66666 million dollars which you would round to 216.7 million dollars.